HUBBLE SPACE TELESCOPE DAILY REPORT #2066 PERIOD COVERED: 0000Z (UTC) 02/23/98 - 0000Z (UTC) 02/24/98 Daily Status Report as of 055/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed FGS 7597 (Long Term Stability of Transfer Scan Mode in FGS-1R) The Fine Guidance Sensors (FGS-1R) was used to make observations of the standard star UPGREN69 in order to study the long term stability of the Transfer Scan Mode. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/1/2/3 7875 (Searching for Old Stellar Populations in Candidate Proto-Spheroidals) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations searching for old stellar populations in candidate proto-spheroid SA57-17731. We have compelling evidence from WF/PC-2 images that these galaxies have optical colors gradients. The near-infrared is necessary, however, to reveal color gradients large enough to separate young and old stellar populations. Only NICMOS imaging provides enough spatial resolution and depth to resolve these gradients. The observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at its nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Four Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD 7636 (STIS Spectroscopic CCD Flatfielding) The Space Telescope Imaging Spectroscope (CCD) was used to make a CCD flat field calibration observation in the spectrographic mode. A bias calibration was also taken. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/CCD 7928 (STIS Spectroscopic CCD Flatfielding -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat field calibration images in the STIS spectroscopic mode. The calibrations were completed as planned, and no problems were reported. 1.10 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxy IR11095-02. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.11 Completed NIC/2 7863 (The Nature of Gamma-Ray Burster Host Galaxies) The Near Infrared Camera (Camera Two) was used to study the nature of a gamma-ray burster in the host galaxy GRB70228. Of the two thousand gamma-ray bursts detected in the last twenty-five years, only two have identified optical counterparts. About two hours into this almost six hour proposal (and as reported in yesterday's report), at 054/082055Z, the NICMOS instrument suspended. At least the final four hours of science was lost. 1.12 Completed STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.13 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.14 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.15 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.16 Completed STIS/MA1 7345 (Origin and Physical Conditions in the Local Lyman-Alpha Forest) The Space Telescope Imaging Spectrograph (MA1) was used to make observations of PG0804+761 in order to study the origin and physical conditions in the local Lyman-alpha forest. A wavelength calibration image was also taken. As reported in 2.1, the acquisition for this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.17 Completed NIC/3 7499 (NICMOS Extragalactic Parallel Grism Survey) The Near Infrared Camera (Camera Three) was to be used to make observations as part of an extragalactic parallel Grism Survey. This proposal is a parallel program using NICMOS's unique capabilities for slitless spectroscopy in the near IR. This proposal did not execute since NICMOS was in suspend. 1.18 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was to be used to make pure parallel spectrographic observations of sky regions when another instrument is prime. This proposal did not execute since NICMOS was in suspend. 1.19 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-2639 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The first iteration of this proposal did not execute since NICMOS was in suspend. The second iteration completed as expected with no anomalies. 1.20 Completed NIC/1/2/3 7854 (The Onset of Axisymmetry in Proto-Planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the axissysmitry of the proto-planetary nebula CIT6. The high-resolution imaging and multicolor polarimetric capabilities of NICMOS comprise a superb combination for studying the processes by which evolved (approximately one solar mass) stars shed their outer layers to form planetary nebulae. About one hour into this almost four hour proposal, at 054/1745Z, the NICMOS was recovered from suspend. 1.21 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Near Infrared Camera (Camera One) was used to make observations of the Seyfert Galaxy NGC-5695. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. The observations were executed as scheduled, and no anomalies were reported. 1.22 Completed NIC/1 7855 (The Evolution of Powerful Radio Galaxies) The Near Infrared Camera (Camera One) was used to make observations of the powerful radio galaxy 4C32.44. Our proposed NICMOS observations combined with our multi-wavelength observing program will finally provide us with the data we need to build a picture of how powerful radio galaxies evolve. The observations were completed as planned, and no porblems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 The acquisition at 054/123402Z defaulted to fine lock backup on FGS-3 only when FGS-1 failed to acquire. HSTAR 6362 was written. The proposal described in 1.16 may have been affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: It was determined that the cause of the NICMOS suspend at 054/082055Z was the detector 3 voltages going out of limits low. A NICMOS memory dump was performed at 054/1158Z per ROP NS-9. Analysis of the dump indicates that this event had the same signature as another event seen on day 62 of last year. NICMOS was successfully recovered from suspend at 054/1745Z. Proposals detailed in 1.11, 1.17, 1.18, 1.19, and 1.20 were affected. A NICMOS memory monitor for a single bit error was established at 054/1811Z. Another memory monitor for the EDAC address was set up at 054/1924Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: DF-224 flight software 10.0 is scheduled to be installed today, beginning at approximately 055/1743Z. Continuation of normal science observations and calibrations. /DMH